# Deep Learning for Contextualized NetFlow-Based Network Intrusion Detection: Methods, Data, Evaluation and Deployment

### v1

@@ 3.4 Unsupervised and Self-Supervised Temporal Representation Learning
Unsupervised and self-supervised methods reduce reliance on labeled attack data by learning temporal representations through reconstruction or contrastive objectives. Masked context reconstruction using transformer architectures improves robustness to label noise and anomalous contamination, yielding AUC gains exceeding 9% on datasets such as UNSW-NB15 and CICIDS-WED (Wang et al., 2023). Self-supervised contrastive learning further enables detection of zero-day attacks and adaptation to concept drift by maximizing agreement between augmented flow sequences (Koukoulis et al., 2025).  Guerra et al. (2025) proposed GraphIDS, combining masked graph autoencoding with contrastive learning to capture both local and global structural patterns while reducing labeled data requirements by up to 80%.
@@ 4 Graph Modeling
Self-Supervised Graph Masked node/edge reconstruction; contrastive graph views. Sparse or weakly labeled attacks. Risk of structural leakage; large unlabeled graphs required. Masking must not reveal full adjacency; cross-dataset testing. Guerra et al. (2025); Caville2022AnomalE; Van Langendonck and others (2024)
@@ 4.1 Homogeneous Graph Neural Networks
Self-supervised learning has also been explored in homogeneous graph settings. Methods such as GraphIDS and Anomal-E employ masked autoencoding and edge reconstruction objectives to reduce reliance on labeled data and improve robustness to noisy annotations (Guerra et al., 2025; Caville et al., 2022). While effective, these approaches require large unlabeled graphs and careful masking strategies to prevent structural information leakage.
@@ 4.2 Heterogeneous Graph Neural Networks
Heterogeneous GNNs extend graph-based NIDS by modeling multiple node and edge types with relation-specific parameters, enabling richer representation of network semantics. Relational Graph Convolutional Networks (R-GCNs) provide a foundational framework for this paradigm by supporting typed message passing across diverse relations (Schlichtkrull et al., 2018).
@@ 4.4 Self-Supervised Graph Representation Learning
Self-supervised graph learning aims to learn transferable representations from unlabeled traffic graphs, reducing dependence on costly labeled data. GraphIDS demonstrates this paradigm by combining masked graph autoencoding with contrastive learning to capture both local and global structural patterns (Guerra et al., 2025). The method achieves strong macro-F1 and PR-AUC scores on NetFlow benchmarks while reducing labeled data requirements by up to 80%.  Caville et al. (2022) proposed Anomal-E, a self-supervised network intrusion detection system based on GNNs that learns effective representations without labeled attack data.
Despite these promising results, self-supervised graph learning for NIDS remains underexplored. Open challenges include the design of effective graph augmentations, prevention of temporal and structural leakage during pre-training, and robustness under cross-dataset evaluation (Caville et al., 2022; Lin et al., 2025). Addressing these issues is critical for deploying self-supervised GNNs in dynamic operational environments.  Van Langendonck and others (2024) introduced PPT-GNN, a practical pre-trained spatio-temporal GNN that enables near real-time predictions and better cross-network generalization.
@@ 7.3 Adaptive Learning and Concept Drift
Operational NIDS must cope with label scarcity and non-stationary traffic distributions. Self-supervised learning reduces reliance on labeled data by leveraging abundant unlabeled traffic through masked prediction and contrastive objectives, enabling robust representation learning with substantially fewer annotations (Guerra et al., 2025; Koukoulis et al., 2025).

### v2

@@ 3.4 Unsupervised and Self-Supervised Temporal Representation Learning
Unsupervised and self-supervised methods reduce reliance on labeled attack data by learning temporal representations through reconstruction or contrastive objectives. Masked context reconstruction using transformer architectures improves robustness to label noise and anomalous contamination, yielding AUC gains exceeding 9% on datasets such as UNSW-NB15 and CICIDS-WED (Wang et al., 2023). Self-supervised contrastive learning further enables detection of zero-day attacks and adaptation to concept drift by maximizing agreement between augmented flow sequences (Koukoulis et al., 2025).  Guerra et al. (2025) proposed GraphIDS, combining graph representation learning and transformer-based autoencoder to capture both local and global structural patterns.
@@ 4 Graph Modeling
Self-Supervised Graph Edge feature reconstruction; random node/edge permutation or masking; negative edge sampling; link prediction. Sparse or weakly labeled attacks. Risk of structural leakage; large unlabeled graphs required. Masking must not reveal full adjacency; cross-dataset testing. Guerra et al. (2025); Caville et al. (2022); Van Langendonck et al. (2025)
@@ 4.1 Homogeneous Graph Neural Networks
Self-supervised learning has also been explored in homogeneous graph settings. Methods such as GraphIDS and Anomal-E employ edge features reconstruction and negative graph (corrupted graph) embedding discrimination objectives to reduce reliance on labeled data and improve robustness to noisy annotations (Guerra et al., 2025; Caville et al., 2022). While effective, these approaches require large unlabeled graphs and careful masking strategies to prevent structural information leakage.
@@ 4.2 Heterogeneous Graph Neural Networks
Heterogeneous GNNs extend graph-based NIDS by modeling multiple node and edge types with relation-specific parameters, enabling richer representation of network semantics. Relational Graph Convolutional Networks (R-GCNs) provide a foundational framework for this paradigm by supporting typed message passing across diverse relations (Schlichtkrull et al., 2018; Van Langendonck et al., 2025).
@@ 4.4 Self-Supervised Graph Representation Learning
Self-supervised graph learning aims to learn transferable representations from unlabeled traffic graphs, reducing dependence on costly labeled data.  Caville et al. (2022) proposed Anomal-E, a self-supervised network intrusion detection system based on GNNs that learns effective representations without labeled attack data. It trains a graph encoder (E-GraphSAGE) using edge‑feature reconstruction and discrimination between real and corrupted graphs, encouraging embeddings that capture both local edge attributes and global structural patterns. For anomaly detection, they utilized four unsupervised algorithms: principal component analysis (PCA), isolation forest (IF), cluster-based local outlier factor (CBLOF), and histogram-based outlier score (HBOS). Each algorithm was trained on edge embeddings derived from the training graph, which served as the input features for learning normal behavioral patterns and identifying anomalies.  (Guerra et al., 2025) introduce GraphIDS, a self-supervised framework for network intrusion detection that jointly optimizes graph representation learning and anomaly detection through an end-to-end architecture combining an inductive graph neural network (E-GraphSAGE) with a Transformer-based masked autoencoder. Across diverse NetFlow benchmarks, GraphIDS acheive up to 99.98% PR-AUC and 99.61% macro F1, exceeding baseline methods by 5–25 percentage points.
Despite these promising results, self-supervised graph learning for NIDS remains underexplored. Open challenges include the design of effective graph augmentations, prevention of temporal and structural leakage during pre-training, and robustness under cross-dataset evaluation (Caville et al., 2022; Lin et al., 2025). Addressing these issues is critical for deploying self-supervised GNNs in dynamic operational environments.  Van Langendonck et al. (2025) introduced PPT-GNN, a practical pre-trained spatio-temporal GNN that enables near real-time predictions and better cross-network generalization. It represents NetFlow traffic as heterogeneous graphs built from short time-based sliding windows, enriched with intra- and inter-window temporal edges. Then it uses a self-supervised link-prediction task for every edge in the graph.
@@ 7.3 Adaptive Learning and Concept Drift
Operational NIDS must cope with label scarcity and non-stationary traffic distributions. Self-supervised learning reduces reliance on labeled data by leveraging abundant unlabeled traffic through masked prediction and contrastive objectives, enabling robust representation learning without relying on labeled data (Guerra et al., 2025; Koukoulis et al., 2025).